Jordan deposited $1,000 into TD Bank at 2.8% annual interest, compounded monthly. Determine the value of her account after 15 years.
Answer by stanbon(75887)

Jordan deposited $1,000 into TD Bank at 2.8% annual interest, compounded monthly. Determine the value of her account after 15 years.
------------------
A(t) = P(1+(r/n))^(nt)
-------
A(15) = 1000(1+(0.028/12))^(12*15)
----
A(15) = 1000(1.0023333..)^(180)
============================
A(15) = $1521.22
======================
Cheers,
Stan H.